OpenStudy (anonymous):

f(x)=2x+5 and g(x)=x-7,solve for f(g(x)) when x=-3

OpenStudy (tester97):

Sorry for the inconvenience of someone not answering your question. Usually the average amount of time for someone to answer it is 5-6 minutes. If you question is not answered within given time, patiently wait a minute or two. If that does not work try tagging a few users that are higher on the "users online" list. I truly hope this was not a bad experience for you. Please feel free to tag me in the future using the "@" symbol like so, @tester97 Have a wonderful day! :)

OpenStudy (neer2890):

f(g(x))=f(x-7)=2(x-7)+5=2x-14+5=2x-9 now put x=-3 f(g(x))=2(-3)-9=-6-9=-15